50% of the nuclei in a radioactive sample decay in one half life. What percent decay in four half lives?
Answer #1

N = No*(1/2)^(t/T1/2)

t = 4T1/2

=> N = No*(1/2)^(4) = No/16 this is the remaining after 4 half lives.

=> % decay = 100*(1 - 1/16) = 93.75 %